### v1

# Follow-Your-Motion: Video Motion Transfer via Efficient Spatial-Temporal Decoupled Finetuning

Abstract
Recently, breakthroughs in the video diffusion transformer have shown remarkable capabilities in diverse motion generations. As for the motion-transfer task, current methods mainly use two-stage Low-Rank Adaptations (LoRAs) finetuning to obtain better performance. However, existing adaptation-based motion transfer still suffers from motion inconsistency and tuning inefficiency when applied to large video diffusion transformers. Naive two-stage LoRA tuning struggles to maintain motion consistency between generated and input videos due to the inherent spatial-temporal coupling in the 3D attention operator. Additionally, they require time-consuming fine-tuning processes in both stages. To tackle these issues, we propose Follow-Your-Motion, an efficient two-stage video motion transfer framework that finetunes a powerful video diffusion transformer to synthesize complex motion. Specifically, we propose a spatial-temporal decoupled LoRA to decouple the attention architecture for spatial appearance and temporal motion processing. During the second training stage, we design the sparse motion sampling and adaptive RoPE to accelerate the tuning speed. To address the lack of a benchmark for this field, we introduce MotionBench, a comprehensive benchmark comprising diverse motion, including creative camera motion, single object motion, multiple object motion, and complex human motion. We show extensive evaluations on MotionBench to verify the superiority of Follow-Your-Motion.
Figure 1: Showcases of our Follow-Your-Motion. Given an input video, Follow-Your-Motion enables generating the video with the same motion, including motion of single or multiple objects, complex poses of humans, and movements of the camera view.
1 Introduction
Motion transfer aims to synthesize new videos with the same motion as the reference video. Unlike video-to-video translation qi2023fatezero ; wu2022tune , motion transfer focuses on extracting and reproducing camera view changes and object motions, rather than preserving the low-level and 2D spatial structure. This task has potential applications in diverse fields, including the film industry, augmented reality, automated advertising systems, and social media content creation.
In recent years, the diffusion model rombach2022high has demonstrated remarkable capabilities in generating high-quality visual content due to its stability when optimized for Gaussian noise removal objectives. In addition, Diffusion Transformer (DiT) has further shown its superiority in scalability of model parameters, computation, and large-scale video data. Utilizing pretrained video diffusion models, various motion transfer techniques are proposed using training-free or fine-tuning paradigms. Training-free motion transfer methods geyer2023tokenflow ; pondaven2025ditflow ; qi2023fatezero ; xiao2024video ; yang2025videograin replace or optimize the generated motion representation with the reference video in inference time, while freezing the parameters of neural networks. e.g., . SMM yatim2024space proposes a new descriptor by averaging video features along spatial dimensions. Motionshop yesiltepe2024motionshop uses the latent updates in the denoising pipeline as Motion Score representation for DiT models. Although these training-free methods avoid the network training cost and generalize in both early UNet and recent DiT models, their generated motion is bottlenecked by the knowledge of the pretrained diffusion model.
Figure 2: Comparison between Follow-Your-Motion and baseline. We finetune the baseline and our method 3,000 steps using Wan2.1 wan2025 . Our method gets better reconstruction and motion preservation.
To generate complex motions beyond the domain of the pretrained diffusion model, tuning-based motion transfer zhao2023motiondirector has emerged to learn the motion in the reference video precisely by optimizing the model parameters. In the training stage, early UNet-based methods (MotionDirector) optimize the parameters of separated temporal layers to learn the input motion. In the inference stage, new videos are synthesized by combining the motion learned in the optimized parameters and prior knowledge in the frozen foundation model. Despite the improvements in motion transfer quality based on UNet models, applying tuning-based motion transfer to modern DiT is still challenging because of its high training cost and spatial-temporal mixed computation in its 3D self-attention blocks. In the context of tuning DiT for motion transfer, a naive baseline is to finetune the whole DiT model using Low-Rank Adaptation (LoRA) in the 3d self-attention layer directly.
Tuning-based motion transfer, including U-Net zhao2023motiondirector and DiT abdal2025dynamic employ a spatial–temporal decoupled tuning paradigm: They first optimize a spatial LoRA on sampled frames to capture spatial structure, then freeze those weights and optimize a temporal LoRA over the full video frames to learn motion dynamics. However, we argue that this two-stage procedure is inherently inefficient. Specifically, the limitations are listed as follows:
  1. (1)
    Motion inconsistency: In the first stage, both spatial and temporal heads in self-attention are tuned on a fixed frame, which leads to spatial-temporal coupling. As shown in the top Fig. 2, the naive baseline tunes both spatial and temporal heads, but both the reconstructed results and motion transfer results fail to follow the dog’s original motion. Therefore, tuning both spatial and temporal heads for appearance preservation is not reasonable.
  2. (2)
    Tuning inefficiency: Motivated by the observation in xi2025sparse , the 3D full attention mechanism in DiT-based models contains attention heads that specialize in either spatial or temporal modeling. However, existing methods tune all attention heads in the spatial/temporal tuning stages, regardless of their roles, leading to inefficient parameter usage and redundancy. Moreover, since 3D VAEs compress temporal sequences into compact latent representations, using all reference frames for motion transfer ignores 3D VAEs’ interpolation ability and introduces redundancy.
To address these challenges, we propose Follow-Your-Motion, an efficient video motion transfer framework. Specifically, to mitigate the motion inconsistency problem, we first apply robust head matching to categorize attention heads into spatial and temporal types. During tuning, only spatial heads are updated in the spatial stage, and only temporal heads in the temporal stage. As shown in Figure 2 (a), our method preserves the motion consistency in both reconstruction and transfer results compared to the naive baseline. To address tuning inefficiency, we introduce sparse motion sampling during the temporal tuning stage to accelerate. Furthermore, we propose adaptive RoPE to enhance the model’s ability to learn motion interpolation, allowing it to effectively capture motion dynamics even from sparse frames. As evidence in Figure 2 (b), our decoupled tuning strategy reduces latency by 1.26×, and with sparse sampling, achieves a 3.89× speed-up compared to tuning with all frames.
With the aforementioned designs, our approach can effectively generate high-quality videos following the motion in the reference video, as illustrated in Figure 1. Additionally, to address the lack of a benchmark in video motion transfer, we introduce a comprehensive benchmark called MotionBench, which consists of 53 various video motions, including single object motion, complex human motion, multi-object motion, and camera motion that showcase a wide range of scenes and styles. Compared with the existing baseline methods, we achieve remarkable results on 8 various evaluation metrics on MotionBench. Our method can facilitate more accurate video motion transfer using powerful DiT. Overall, our contributions can be summarized as follows:
  • We propose Follow-Your-Motion, a two-stage video motion transfer framework that tunes the powerful video diffusion transformer to synthesize video clips with complex motion.
  • We analyze motion inconsistency and tuning inefficiency in DiT-based video motion transfer. To ensure motion consistency, we separately tune spatial and temporal LoRA heads for appearance and motion, respectively. To address tuning inefficiency, we introduce adaptive RoPE and sparse motion sampling to accelerate training.
  • To validate the effectiveness of our methods, we construct a benchmark MotionBench. We perform extensive experiments and user studies to evaluate our approach, which shows our method achieves state-of-the-art performance.
2 Related Work
Text-to-video generation. Text-to-video generation aims to produce realistic videos that precisely match the spatial visuals and temporal dynamics described in the input prompt. To generate the complicated motion in the videos, diffusion-based video generation models guoanimatediff ; ma2024follow ; ma2022visual ; zhang2024follow ; zhao2023controlvideo ; zhang2025magiccolor ; ma2024followyourclick ; zhu2025multibooth ; liu2025avatarartist ; yan2025eedit ; wan2024unipaint ; chen2024follow ; wang2024taming are proposed to synthesize consistent results using a pretrained image diffusion model. Previous works chen2024m ; chen2023attentive ; guo2023animatediff ; He2022LatentVD ; wang2023modelscope ; ma2024followyouremoji ; xiong2025enhancing ; wang2024cove ; feng2025dit4edit ; zhu2024instantswap design the temporal module of UNet to generate consistent results. Recently, the emergence of Diffusion Transformer-based methods for text-to-video generation has exhibited superior performance in quality and consistency. These powerful scaling transformers, including Sora Liu2024SoraAR , CogVideoX Yang2024CogVideoXTD , EasyAnimate xu2024easyanimate , HunyuanVideo kong2024hunyuanvideo , and Wan2.1 wan2025 , enable generating more realistic video clips from given detailed prompts, paving the way for various downstream video generation tasks.
Video Motion transfer. Motion transfer involves an important demand: creating a novel video and maintaining the motion from the reference one. Some methods leverage the explicit control signal ma2023magicstick ; ma2024follow ; xing2024make ; xing2024dynamicrafter ; zhang2025framepainter to achieve motion transfer from the reference video. However, these methods rely on a huge control signal dataset and cost large computational resources. Thanks for the powerful pretrained text-to-video generation model, the researchers pay attention to motion transfer using implicit control, including training-free or tuning-based paradigm. For training-free methods hu2024motionmaster ; pondaven2025ditflow ; yesiltepe2024motionshop , they extract a motion embedding in the inference stage and use the gradient to guide optimization. However, these methods fail to transfer the complex motion. For tuning-based methods, they jeong2024dreammotion ; zhao2023motiondirector always fine-tune model parameters to utilize different attention for temporal and spatial information. Current works jeong2024dreammotion ; ren2024customize ; zhao2023motiondirector employ the dual-path LoRA structure to separate motion and appearance. However, these methods are developed on the UNet-based pretrained model chai2023stablevideo , making them unsuitable for DiTs. In contrast, our proposed method is the first one-shot DiT-based motion transfer framework. Using the video diffusion transfer as the foundation model, our method extends the boundary of motion transfer performance.
3 Method
We target video motion transfer from the input source video to the new synthesized one by optimizing generative models to learn the object motions and camera pose changes in the source video. As proposed in previous work abdal2025dynamic ; zhao2023motiondirector , a naive baseline (Fig. 2) of motion transfer is to first optimize the LoRA weight for sampled video frames as a text-to-image task, where the LoRA obtained in the first stage is called spatial LoRA ($\Delta W_{s}$). In the second stage, consecutive multiple frames are treated as ground truth to optimize the temporal LoRA ($\Delta W_{t}$), while the spatial LoRA is loaded and kept frozen. During the inference stage, only spatial lora ($\Delta W_{t}$) is loaded to transfer the video motion in the source video while presenting the appearance bias leakage. Although this strategy can generate results for the UNet-based diffusion model, the computation cost for the DiT-based video diffusion model is still too expensive. As shown in Fig. 2, such naive LoRA tuning cannot reproduce the motion in the input video after tuning 3000 steps (3042 seconds on a single H20 GPU). Previous naive LoRA tuning faces two main challenges. (1) Recent Video DiT models leverage 3D attention block without explicit temporal blocks, which makes it difficult to disentangle temporal parameters and fine-tuning LoRA on whole attention parameters results in larger parameter number (e.g., 29.5 M for naive LoRA) (2) Finetuning on multiframe videos increases token sequence length (e.g., 24276 tokens for 81 frames) and computation cost.
To address these challenges, in Sec 3.1, we first present an attention head classification strategy to decouple the spatial and temporal model parameters by analyzing the sparsity of attention in the pretrained video diffusion transformer. Then we introduce an efficient and economical tuning framework (Sec. 3.2) for learning spatial appearance and temporal motion in the source video. Note that we use WAN as our pretrained backbone in our experiments, but our method can generalize to other video DiT models.
Figure 3: Overview of our methods. Stage 1: We first classify the attention heads using a pseudo spatial attention map. Stage 2: After attention classification, we first tune the spatial LoRA using a random frame in the video. Stage 3: After finishing spatial LoRA tuning, we load the spatial LoRA weight and conduct temporal tuning using sparse motion sampling and adaptive RoPE.
3.1 Stage 1: Spatial-temporal attention classification
The pretrained video DiT model Wan utilizes unified 3D attention instead of separated spatial and temporal attention, which brings challenges to motion information decoupling pondaven2025ditflow , training efficiency, and storage cost pondaven2025ditflow . Inspired by evidence in previous work xi2025sparse , we leverage the inherent sparsity in 3D Full Attention of video DiT to decouple the parameters for temporal motion and spatial appearance.
  Input: $Q,K,\in\mathbb{R}^{H\times S\times D}$: query and key where sequence $S=F\times H\times W$
1
Output: Closest head type: $t_{\text{head}}$
2
[…]
6 $\triangleright$ Get attention maps of input data:$[\text{head},S,S]$
[…]
8 $\triangleright$ Calculate similarity metrics
[…]
10 $\text{Sim}_{t}\leftarrow\left\|M_{\text{input}}\odot M_{\text{t}}\right\|.\text{mean}(\text{dim}=(1,2))$
[…]
$t_{\text{head}}$leftarrow ($\text{Sim}_{s}<\alpha\text{Sim}_{t}$)
Algorithm 1 Dual attention decoupling
Dual attention decoupling.
As shown in Alg. 1, our method classifies attention heads in Wan into temporal or spatial types. We take query and key tokens $Q,K\in\mathbb{R}^{H\times S\times D}$ as input, where $H$ is the number of heads, $S$ is the sequence length, and $D$ is the feature dimension.
We prepare pseudo ground truths: for spatial attention map, $M_{\text{spatial}}[i,j]=1$ if points $(i,j)$ are near the main diagonal ( within a predefined range), otherwise 00; for temporal attention map, $M_{\text{temporal}}[i,j]=1$ if points $(i,j)$ are near diagonals parallel to the main diagonal (identical spatial positions in different frames), otherwise 00.
We compute the cosine similarity $\text{Sim}_{s}$ between the input attention map $M_{\text{input}}$ and $M_{\text{spatial}}$, and $\text{Sim}_{t}$ between $M_{\text{input}}$ and $M_{\text{temporal}}$. A head is classified as temporal if $\text{Sim}_{s}<\alpha\cdot\text{Sim}_{t}$, where $\alpha=1.3$(empirically set to balance the number of spatial and temporal heads).
Dual attention fusion. Then, we rearrange the channels of the linear layers in full 3D attention $q,k,v,o$ to two parallel branches for temporal attention and spatial attention. The forward algorithm of a single rearranged block is shown in Alg. 2. Given input sequence $x$, we concatenate the features from the temporal and spatial branches along the channel dimensions to get the tokens of query $Q$, key $K$, and value $V$. After applying rotary position embedding and scaled dot product attention, feature $x$ is split along the channel dimension, and fed to $\mathrm{o}_{temp}$ and $\mathrm{o}_{spat}$. Finally, the summed feature is returned at the end of the attention block.
3.2 Stages 2&3: Spatial-temporal decoupled tuning
Spatial LoRA tuning. As the parameters of the attention block are decoupled in the previous stage, we can use the spatial and temporal branches in two stages to learn the appearance and motion in the reference videos, respectively. Following previous work abdal2025dynamic ; zhao2023motiondirector , we first inject LoRAs $\theta_{spat}$ into the spatial heads branch $(q_{spat},k_{spat},v_{spat},o_{spat})$ to learn the spatial appearance in stage 1. We randomly sample a batch of frames $x_{i}$ from index $\{0,1,2,...,F-1\}$, and optimize the spatial LoRA $\theta_{spat}$ as a text-to-image model using the training loss:
$\mathcal{L}_{spat}=E_{x_{1},x_{0}\sim N(0,I),i\sim U(0,F)}\left\|v_{t,i}-v_{\theta_{spat}}\left(x_{t,i},t,p\right)\right\|_{2}^{2}.$ (1)
Temporal LoRA tuning. Once the spatial LoRA $\theta_{spat}$ gets converged, we freeze $\theta_{spat}$ in the model, and continue to finetune temporal LoRA parameters $\theta_{spat}$ of temporal heads branch $(q_{temp},k_{temp},v_{temp},o_{temp})$. Since the Wan wan2025 is pretrained on a large frame number $F=81$, fine-tuning on the original number $F=81$ costs too expensive computation (Fig. 2). To alleviate the high computation requirement for videos, we propose the sparse motion sampling, which finetune our spatial LoRA ${\theta_{spat}}$ on a sampled video with fewer frame number $F_{samp}=17$ and then infer with the original frame number. While recent transformer models apply Rotary Positional Embedding (RoPE) vaswani2017attention to encode the relative position dependency according to the frame index, sampling frames from $F$ to $F_{samp}$ breaks the original dependency and thus deteriorates the motion quality. Motivated by previous text-to-image DiT models kong2024hunyuanvideo ; yang2024cogvideox , we propose the adaptive RoPE, a centralized scaling positional encoding along the frame index to align the position range with different total frame numbers. For each frame with temporal index $i\in[0,1,...,{F_{samp}-1}]$, its temporal positional embedding is assigned as:
$\text{PE}_{x_{i}}=f(\frac{F}{2}+\frac{F}{F_{samp}}(i-\frac{F_{samp}}{2})),$ (2)
which ensures that videos with less frame number $F_{samp}$ have the same input range $[0,F]$ for the embedding function $f$, as the pertaining stage of video DiT.
Figure 4: Illustration of sparse motion sampling and adaptive RoPE. The adaptive RoPE is utilized to represent frame position in the video.
Input:
$x\in\mathbb{R}^{H\times S\times D}$: input sequence
$f_{\mathrm{regs}}$: positional frequencies
$d_{\mathrm{temp}}$: temporal dimension size
Output: Fused output $y\in\mathbb{R}^{H\times S\times D}$
1 $\triangleright$ Channel concatenate, and normalize
2 $Q\leftarrow\mathrm{Norm}([q_{\mathrm{temp}}(x)\|q_{\mathrm{spat}}(x)])$
3 $K\leftarrow\mathrm{Norm}([k_{\mathrm{temp}}(x)\|k_{\mathrm{spat}}(x)])$
4 $V\leftarrow[v_{\mathrm{temp}}(x)\|v_{\mathrm{spat}}(x)]$
5 $\triangleright$ Rotary Position Embeddings for Q, K
6 $\widetilde{Q}\leftarrow\mathrm{RoPE}(Q,f_{\mathrm{regs}},H)$
7 $\widetilde{K}\leftarrow\mathrm{RoPE}(K,f_{\mathrm{regs}},H)$
8 $\triangleright$ Multi-Head Attention
9 $x\leftarrow\mathrm{Attention}(\widetilde{Q},\,\widetilde{K},\,V;\,H)$
10 $\triangleright$ Dual Output Projection Fusion
11 $y\leftarrow o_{\mathrm{temp}}(x[:d_{\mathrm{temp}}])+o_{\mathrm{spat}}(x[d_{\mathrm{temp}}:])$
return $y$
Algorithm 2 Dual Attention Fusion
To further decouple the temporal motion from spatial appearance, we further introduce a motion loss ling2024motionclone by eliminating the appearance and focusing on the changes in the temporal dimensions. We first define the motion latent $\hat{v}$ for each frame $i$ as the difference between the current frame latent $v_{i}$ and the previous frame latent $v_{i-1}$ as: $\hat{v}_{i}=v_{t,i}-v_{t,i-1}.$
Then, we define the motion loss following zhao2023motiondirector as the negative cosine similarity between the ground truth motion latent and predicted motion latent:
$\mathcal{L}_{Motion}=E_{x_{1},x_{0}\sim N(0,I),i\sim U(0,F)}[1-\text{CosineSim}(\hat{v}_{t,i},\hat{v}_{\theta}\left(x_{t,i},t,p\right))].$ (3)
Finally, the total loss for temporal LoRAs is the combination of general video denoising loss and motion loss as $\mathcal{L}_{temp}$ = $\mathcal{L}_{\text{{video\_denoise}}}$ + $\mathcal{L}_{Motion}$.
Figure 5: Gallery of our proposed methods. Given a reference video, our Follow-Your-Motion capability of generating a high-quality video clip with the same motion, including single object motion, multiple object motion, complex human motion, and camera motion.
4 Experiments
4.1 Implementation details
In our experiment, we employ the open-sourced video generation model WAN-2.1 wan2025 as the base text-to-video generation model. The LoRA ranks are 128 in both stages. We first randomly select a single frame and take about 3,000 steps for spatial appearance learning. The AdamW loshchilov2017decoupled optimizer is utilized, and the learning rate is $1\times 10^{-5}$. The spatial weight decay is $0.1$. During the second tuning stage, we freeze the spatial head LoRA and only train the temporal head LoRA for 2000 steps with learning rate $1\times 10^{-5}$ and weight decay $0.99$. For sparse motion sampling, we set the sampling stride to 5. The input videos are fed into the model as $512\times 512$. The sampled frame number is 16 in the second stage. In the inference stage, we leverage the flow matching scheduler lipman2023flow with a sampling step of 30, and a text-guidance ratio of 7.0. The LoRA weights are set as 0.5. More details and evaluation metrics can be found in the supplementary materials.
Figure 6: Qualitative comparison with baselines. We perform the visual comparison with six baselines using various kinds of motions. Our method obtains better performance in various motions.
4.2 MotionBench
In order to address the lack of a benchmark in video motion transfer, we introduce MotionBench, a comprehensive benchmark to evaluate the ability of current motion transfer approaches. In detail, we collect 53 videos from four aspects, including 1). camera motion, 2). single object motion, 3). multiple object motion, and 4). complex human motion. Single object motion sequences focus on diverse motion patterns from a single subject. Multiple object motion involves the consistency of spatial relationships between different instances. Camera motion evaluates viewpoint changes through both simple camera trajectories (zoom, tilt, pan) and complex camera operations. Here, single/multi-object refers to general objects and animals, while human motion contains more non-rigid deformations, so we treat it separately. The 30% videos in our benchmark are generated by text-to-video generation models wan2025 , and other videos are obtained from publicly licensed video websites. We use the GPT4o openai2024gpt4o to get the video captions. In MotionBench, each video is approximately 5 seconds long with 150 frames. MotionBench provides a standardized evaluation protocol across diverse motion categories, enabling systematic assessment and comparison of motion transfer methods.
4.3 Comparison with baselines
Qualitative comparison. We compare our approach with previous video motion transfer methods visually, including state-of-the-art video motion transfer methods: MOFT xiao2024video , MotionInversion wang2024motioninversion , MotionClone ling2024motionclone , SMM yatim2024space , MotionDirector zhao2023motiondirector , DiTFlow pondaven2024videomotion . We exclude Motionshop yesiltepe2024motionshop and MotionCrafter zhang2023motioncrafter . from our comparisons as no public release exists. Our experimental results exhibit Follow-Your-Motion better performance and versatility across diverse motion transfer scenarios. As illustrated in Fig. 6, in single object motion cases (first column), we find that the previous works fail to follow source motion. In contrast, our approaches effectively transform the motion from the source video into the target object, maintaining a consistent motion pattern. For multi-object cases, MotionDirector zhao2023motiondirector and SMM yatim2024space have the challenge of handling multi-object interaction motion. Our method enables generating videos with aligned movement patterns, preserving the spatial relationships between moving subjects. Additionally, we provide a visual comparison of complex camera motion. The visual results demonstrate the superiority of our methods in camera motion transfer capabilities.
Quantitative comparison. We compare our method with state-of-the-art video motion transfer on our MotionBench, and the results are shown in Tab. 1. Due to the limited video length of previous works, all evaluations are performed in 32 frames at a resolution of $512\times 512$. Here, we classify the SOTA methods as two classes, training-free or tuning-based, according to whether they use spatial/temporal LoRA to optimize complex motion patterns. (a) Time: Thanks to sparse motion sampling, Follow-Your-Motion is the fastest tuning-based method. Moreover, our running time is on par with training-free approaches while delivering superior performance. (b) Motion Fidelity: Following  yatim2024space , motion fidelity is applied to evaluate tracklet similarity between reference and output videos. (c) Temporal Consistency: We evaluate the average frame-to-frame coherence using CLIP radford2021learning feature similarity among consecutive video frames. (d)Text similarity: We use CLIP to extract target video features and compute the average cosine similarity between the input prompt and all video frames. (f) User study: Since automatic metrics often fail to reflect real preferences, we invited 20 volunteers to rank methods on MotionBench across four aspects—motion preservation, appearance diversity, text alignment, and overall quality—on a 1 (best) to 7 scale. The average rank per method (lower is better) is shown in Tab. 1. Our method achieves the top result in both automatic metrics and human preference.
Table 1: Comparison with state-of-the-art video motion transfer methods. Red and Blue denote the best and second best results, respectively. User study scores report the average rank (1=best, 7=worst) for all the methods; lower is better.
Method Quantitative Metrics User Study
Text Sim.↑ Motion Fid.↑ Temp. Cons.↑ Time (s)↓ Motion Pres.↓ App.↓ Text Align.↓ Overall↓
Training-Free Methods
MOFT xiao2024video 0.283 0.786 0.928 1241 6.883 6.631 5.894 6.639
MotionClone ling2024motionclone 0.298 0.825 0.906 1028 6.283 5.874 6.642 4.192
SMM yatim2024space 0.276 0.927 0.924 785 4.350 5.086 4.205 5.883
DiTFlow pondaven2025ditflow 0.371 0.813 0.937 714 3.326 2.417 2.215 3.284
Tuning-Based Methods
MotionInversion jeong2024vmc 0.301 0.838 0.764 2359 5.417 3.295 5.117 5.074
MotionDirector zhao2023motiondirector 0.285 0.903 0.943 3022 2.217 4.208 3.298 2.216
Ours 0.375 0.976 0.971 781 1.123 1.335 1.174 1.132
4.4 Ablation study
In this section, we conduct a systematic ablation study to isolate and quantify the contribution of each key component in our framework. The qualitative and quantitative ablation study results are shown in Fig. 7 and Tab. 4.4, respectively.
Table 2: Quantitative ablation. Red and Blue denote best, 2nd.
Method Text Sim.$\uparrow$ Motion Fid.$\uparrow$ Temp. Cons.$\uparrow$ Time(s)$\downarrow$
w/o STD LoRa 0.365 0.542 0.842 983
w/o Adaptive RoPE 0.368 0.658 0.814 788
w/o Sparse Sampling 0.369 0.979 0.968 2053
Ours 0.372 0.976 0.971 781
Figure 7: Ablation study about proposed modules. We remove the proposed modules to evaluate their effectiveness. “STD” means spatial–temporal decoupled LoRA.
Effectiveness of spatial–temporal decoupled LoRA. As shown in the second row of Fig. 7 and the “w/o STD LoRA” ablation in Tab. 4.4, the naive baseline jointly tunes without separating spatial and temporal attention heads, failing to decouple the dog’s appearance and causing the edited tiger to look unnaturally black. In contrast, our decoupled LoRA preserves motion while effectively modifying appearance, as evidenced by the improved text similarity in Tab. 4.4.
Effectiveness of adaptive RoPE. Thanks to our adaptive RoPE design, the model can precisely infer each sampled frame’s original index under sparse motion sampling, ensuring the edited motion remains aligned with the source. Without adaptive RoPE, the tiger’s motion becomes disordered and fails to match the original video dynamics. In Tab. 4.4, an improvement of about 48.3% over motion fidelity, quantitatively confirms the benefit of our adaptive RoPE.
Effectiveness of sparse motion sampling. By employing sparse motion sampling in the temporal tuning phase, we reduce the tuning time from 2053s to 781s—an approximately 2.6$\times$ speed-up. Note that in the “w/o sparse sampling” setting, we still apply adaptive RoPE but tune on all video frames, resulting in identical motion fidelity (0.979 vs. 0.976) while incurring the higher time cost.
5 Conclusion
In this paper, we propose Follow-Your-Motion, a two-stage video motion transfer framework that tunes the video diffusion transformer to synthesize video clips with complex motion. In detail, we analyze the motion inefficiency and tuning inefficiency in DiT-based video motion transfer. Through the proposed efficient spatial-temporal decoupled LoRA, we achieve better motion consistency. To address the tuning inefficiency, we introduce adaptive RoPE and sparse motion sampling to accelerate training. Extensive experimental results demonstrate the effectiveness of our method, which outperforms a wide range of previous methods, achieving state-of-the-art video motion transfer quality.

### v2

@@
 […]
-6 $\triangleright$ Get attention maps of input data:$[\text{head},S,S]$
+6 $\triangleright$ Get attention maps of input data:[head,S,S][$\text${head},S,S]
 […]
@@
 10 $\text{Sim}_{t}\leftarrow\left\|M_{\text{input}}\odot M_{\text{t}}\right\|.\text{mean}(\text{dim}=(1,2))$
-[…]
-$t_{\text{head}}$leftarrow ($\text{Sim}_{s}<\alpha\text{Sim}_{t}$)
+11 $\triangleright$ Classify head type: Boolean tensor $\text{head}]$
+$t_{\text{head}}$ $\leftarrow$ ($\text{Sim}_{s}<\alpha\text{Sim}_{t}$)
 Algorithm 1 Dual attention decoupling
@@
 We take query and key tokens $Q,K\in\mathbb{R}^{H\times S\times D}$ as input, where $H$ is the number of heads, $S$ is the sequence length, and $D$ is the feature dimension.
-We prepare pseudo ground truths: for spatial attention map, $M_{\text{spatial}}[i,j]=1$ if points $(i,j)$ are near the main diagonal ( within a predefined range), otherwise 00; for temporal attention map, $M_{\text{temporal}}[i,j]=1$ if points $(i,j)$ are near diagonals parallel to the main diagonal (identical spatial positions in different frames), otherwise 00.
+We prepare pseudo ground truths: for spatial attention map, $M_{\text{spatial}}[i,j]=1$ if points $(i,j)$ are near the main diagonal ( within a predefined range), otherwise 0; for temporal attention map, $M_{\text{temporal}}[i,j]=1$ if points $(i,j)$ are near diagonals parallel to the main diagonal (identical spatial positions in different frames), otherwise 0.
 We compute the cosine similarity $\text{Sim}_{s}$ between the input attention map $M_{\text{input}}$ and $M_{\text{spatial}}$, and $\text{Sim}_{t}$ between $M_{\text{input}}$ and $M_{\text{temporal}}$.
@@
 Table 2: Quantitative ablation. Red and Blue denote best, 2nd.
-Method Text Sim.$\uparrow$ Motion Fid.$\uparrow$ Temp. Cons.$\uparrow$ Time(s)$\downarrow$
+[…]
 w/o STD LoRa 0.365 0.542 0.842 983
@@
 4.4, an improvement of about 48.3% over motion fidelity, quantitatively confirms the benefit of our adaptive RoPE.
-Effectiveness of sparse motion sampling.
-By employing sparse motion sampling in the temporal tuning phase, we reduce the tuning time from 2053s to 781s—an approximately 2.6$\times$ speed-up.
-Note that in the “w/o sparse sampling” setting, we still apply adaptive RoPE but tune on all video frames, resulting in identical motion fidelity (0.979 vs.
-0.976) while incurring the higher time cost.
+[…]
 5 Conclusion